OpenStudy (anonymous):

An engineer in a locomotive sees a car stuck on the track at a railroad crossing in front of the train. When the engineer first sees the car, the locomotive is 150 m from the crossing and its speed is 16 m/s. If the engineer’s reaction time is 0.75 s, what should be the magnitude of the mini- mum deceleration to avoid an accident?

OpenStudy (shane_b):

Ok, let's start by writing everything out: d=150m v=16m/s reaction time = 0.75s First figure out how much closer the train got in the 0.75s before the engineer started braking: \[(0.75s)(16m/s)=12m\]Therefore, he needs to stop in \[150m-12m=138m\] Now just use the kinematic equation:\[V_f^2=V_i^2+2ad\]Plug in values:\[(0m/s)^2=(16m/s)^2+2a(138m)\]Solving for a you should end up with a braking acceleration of 0.93m/s^2. This could be negative depending on your coordinate system.

OpenStudy (anonymous):

Thank you so much! I got it down to the basics but I didn't know which equation to use,

OpenStudy (shane_b):

No problem, good luck :)